# A robust parameterized enhanced shift-splitting preconditioner for three-by-three block saddle point problems

### 27 Feb 2024 v1

@@ Abstract
This paper proposes a new parameterized enhanced shift-splitting (PESS) preconditioner to solve the three-by-three block saddle point problem (SPP). In addition, necessary and sufficient criteria are established for the convergence of the proposed PESS iterative process for any random initial guess. Furthermore, we meticulously investigate the spectral bounds of the PESS preconditioned matrix. Moreover, empirical investigation has been performed for the sensitivity analysis of the system, revealing the robustness of the proposed PESS preconditioner. Numerical experiments are carried out to demonstrate the enhanced efficiency and robustness of the proposed PESS preconditioner compared to the existing block diagonal and shift-splitting preconditioners.
Keywords. Krylov subspace iterative process $\cdot$ Preconditioner $\cdot$ Shift-splitting $\cdot$ Saddle point problem $\cdot$ Robust AMS subject classification. 65F08 $\cdot$ 65F10 $\cdot$ 65F50
@@ 1 Introduction
where $\mathcal{A}=\left[\begin{matrix}A&B^{T}&0\\ -B&0&-C^{T}\\ 0&C&0\end{matrix}\right],\,{\bf u}=\left[\begin{matrix}x\\ y\\ z\end{matrix}\right],\,{\bf d}=\left[\begin{matrix}f\\ g\\ h\end{matrix}\right],$ and $A\in{\mathbb{R}}^{n\times n}$ is a symmetric positive definite (SPD) matrix, $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ are the full row rank matrices. Here, $B^{T}$ stands for the transpose of $B,$ and $f\in{\mathbb{R}}^{n},\,g\in{\mathbb{R}}^{m}$ and $h\in{\mathbb{R}}^{p}$ are known vectors. Under the stated assumptions on block matrices A, B, and C, $\mathcal{A}$ is nonsingular, which ensures the existence of a unique solution for the system (1.1). Due to the growing popularity of three-by-three block SPPs in a variety of applications, it is essential to determine the efficient and robust numerical solution of three-by-three block SPPs (1.1). However, the indefiniteness and poor spectral properties of $\mathcal{A}$ makes challenging to determine the solution of the system (1.1). Furthermore, $\mathcal{A}$ is generally sparse and large; therefore, iterative processes become superior to direct methods. The Krylov subspace iterative process is preferable among other iterative processes investigated recently in the literature due to their minimal storage requirement and feasible implementation [29]. Slow convergence of the Krylov subspace iterative process is a major drawback for system of linear equations of large dimensions. Moreover, the saddle point matrix $\mathcal{A}$ can be very sensitive, which leads to a significant slowdown in the solution algorithm. Therefore, it is important to develop novel, efficient, and robust preconditioners for the fast convergence of the Krylov subspace iterative process, which can handle the sensitivity of the problem (1.1) when the matrix $\mathcal{A}$ is ill-conditioned. Consider the following partitioning of $\mathcal{A}:$
$\left[\begin{array}[]{cc|c}A&B^{T}&0\\ -B&0&-C^{T}\\ \hline\cr 0&C&\hfil 0\hfil\\ \end{array}\right],\qquad\left[\begin{array}[]{c|cc}A&B^{T}&0\\ \hline\cr-B&0&-C^{T}\\ 0&C&0\\ \end{array}\right],$ (1.8)
@@
then the linear system (1.1) can be identified as the standard two-by-two block SPP [5]. In the last decades, colossal attention has been given to the development of numerical solution methods for standard two-by-two block SPP, such as Uzawa methods [7, 13, 2], Krylov subspace iterative process [17, 26], null space methods [15, 25]. However, the properties of the submatrices in (1.8) are different from the standard two-by-two block SPP. Notice that the leading blocks are skew-symmetric and positive semi-definite in the first partitioning. The second partitioning highlights that the $(1,2)$ block is rank deficient, and the $(2,2)$ block is skew-symmetric. In contrast, in the standard two-by-two block SPP, $(1,2)$ block is full row rank, and the $(2,2)$ block is generally symmetric positive semidefinite. Therefore, the existing literature for solving the two-by-two SPPs (see [5]) may not be applied to solve (1.1) directly, and it is essential to develop new preconditioners for three-by-three block SPP that exploit the specific structure of the saddle point matrix $\mathcal{A}$, which is sensitive in nature.
@@
${\mathscr{P}}_{BD}=\left[\begin{matrix}A&0&0\\ 0&S&0\\ 0&0&CS^{-1}C^{T}\end{matrix}\right]\quad\mbox{and}\quad{\mathscr{P}}_{IBD}=\left[\begin{matrix}\hat{A}&0&0\\ 0&\hat{S}&0\\ 0&0&C\hat{S}^{-1}C^{T}\end{matrix}\right],$ (1.15)
where $S=BA^{-1}B^{T},$ $\hat{A}$ and $\hat{S}$ are SPD approximations of $A$ and $S,$ respectively. Although the spectrum of the preconditioned matrices ${\mathscr{P}}^{-1}_{BD}\mathcal{A}$ and ${\mathscr{P}}^{-1}_{IBD}\mathcal{A}$ have good clustering properties, these preconditioners have certain shortcomings, such as they are time-consuming, expensive, require excessive number of iterations, and condition numbers are very large. For more details, see the paper [18]. To overcome these drawbacks of BD preconditioners, researchers have started exploring shift-splitting (SS) preconditioners, which was initially developed for a non-Hermitian system of linear equations by Bai et al. [3] and later for two-by-two block SPPs [10, 23, 11]. Recently, Cao [9] enhanced this idea and introduced the following SS preconditioner ${\mathscr{P}}_{SS}$ and relaxed SS (RSS) preconditioner ${\mathscr{P}}_{RSS}$ for the three-by-three block SPP (1.1)
${\mathscr{P}}_{SS}=\frac{1}{2}\left[\begin{matrix}\alpha I+A&B^{T}&0\\ -B&\alpha I&-C^{T}\\ 0&C&\alpha I\end{matrix}\right],\,\,\mbox{and}\,\,{\mathscr{P}}_{RSS}=\frac{1}{2}\left[\begin{matrix}A&B^{T}&0\\ -B&\alpha I&-C^{T}\\ 0&C&\alpha I\end{matrix}\right],$ (1.16)
where $\alpha>0$ and verified unconditionally convergence of the associated SS iterative process. Here, $I$ stands for the identity matrix of the appropriate dimension. Wang and Zhang [27] generalized the SS preconditioner by introducing a new parameter $\beta>0$ in the $(3,3).$ Further, Yin et al. [30] developed an extensive generalized SS (EGSS) preconditioner
@@
for the three-by-three block SPP (1.1), where $\alpha,\beta,\gamma>0$ and $P,Q,W$ are SPD matrices and investigated its convergent properties. There are a number of other preconditioners and iterative processes to solve the three-by-three block SPP (1.1) has been studied recently in the literature  [19, 1, 24, 28, 4]. Despite exhibiting favorable performance of SS, RSS and EGSS preconditioners, they do not outperform BD and IBD preconditioners. Therefore, to improve the convergence speed and efficiency of the preconditioners ${\mathscr{P}}_{SS},{\mathscr{P}}_{RSS}$ and ${\mathscr{P}}_{EGSS},$ this paper presents a parameterized enhanced shift-splitting (PESS) preconditioner by introducing a parameter $s>0$ for there-by-three block SPPs. As per the direct correlation between the rate of convergence in Krylov subspace iterative processes and the spectral properties of the preconditioned matrix, we perform in-depth spectral distribution of PESS preconditioned matrix. We summarize the following main contributions of this paper:
  • We propose the PESS iterative process and associated PESS preconditioner to solve the three-by-three block SPP (1.1). The proposed preconditioner generalizes the other SS preconditioners studied in the literature see, for example, [9, 27, 30].
@@
  • We empirically show that our proposed PESS preconditioner significantly reduces the condition number of the ill-conditioned saddle point matrix $\mathcal{A}$. Thereby establishing an efficiently solvable, well-conditioned system.
  • Numerical experiments show that the proposed PESS preconditioner outperforms the existing BD and SS preconditioners. Moreover, we perform sensitivity analysis by perturbing the ill-conditioned matrix $\mathcal{A}$ in (1.1), demonstrating the robustness of the proposed preconditioner.
The structure of this article is as follows. In Section 2, we propose the PESS iterative process and the associated PESS preconditioner. Section 3 is devoted to investigating the convergence of the PESS iterative process. In Section 4, we investigate the spectral distribution of the PESS preconditioned matrix. In Section 5, we conduct numerical experiments to illustrate the computational efficiency and robustness of the proposed preconditioner. In the end, a brief concluding remark is provided in Section 6. Notation. The following notations are adopted throughout the paper. For any matrix $A\in{\mathbb{R}}^{n\times n},$ $A^{T},\sigma(A)$ and $\vartheta(A)$ denotes transpose, spectrum and spectral radius of $A,$ respectively. For any $x\in\mathbb{C}^{n},$ $x^{*}$ denotes its conjugate transpose. The symbol $\mathrm{diag}(A,B)$ represents a block diagonal matrix with diagonal blocks $A$ and $B.$ The real and imaginary parts of any complex number $z$ are represented by $\Re(z)$ and $\Im(z)$, respectively. For any matrix $A\in{\mathbb{R}}^{n\times n},$ $\lambda_{\max}(A)$ and $\lambda_{\min}(A)$ denotes the maximum and minimum eigenvalue of $A,$ respectively. The notation $A>0$ signifies that $A$ is SPD, and $A>B$ identifies $A-B>0.$ Abbreviations used in the paper are as follows:
SPP: saddle point problem, SPD: symmetric positive definite, BD: block diagonal, IBD: inexact block diagonal, SS: shift-splitting, RSS: relaxed shift-splitting, EGSS: extensive generalized shift-spliting, PESS: parameterized enhanced shift-splitting, GMRES: generalized minimum residual process.
@@ 2 The parameterized enhanced shift-splitting (PESS) iterative process and the preconditioner
Let $s>0$ be a real number. Then, we write the matrix $\mathcal{A}$ in the form
$\mathcal{A}=(\Sigma+s\mathcal{A})-(\Sigma-(1-s)\mathcal{A}):={\mathscr{P}}_{PESS}-\mathcal{Q}_{PESS},\,\mbox{where}$ (2.1)
${\mathscr{P}}_{PESS}=\left[\begin{matrix}\Lambda_{1}+sA&sB^{T}&0\\ -sB&\Lambda_{2}&-sC^{T}\\ 0&sC&\Lambda_{3}\end{matrix}\right],\mathcal{Q}_{PESS}:=\left[\begin{matrix}\Lambda_{1}-(1-s)A&-(1-s)B^{T}&0\\ (1-s)B&\Lambda_{2}&(1-s)C^{T}\\ 0&-(1-s)C&\Lambda_{3}\end{matrix}\right],$
$\Sigma=\mathrm{diag}(\Lambda_{1},\Lambda_{2},\Lambda_{3}),$ and $\Lambda_{1},\Lambda_{2},\Lambda_{3}$ are SPD matrices. The matrix ${\mathscr{P}}_{PESS}$ is nonsingular for $s>0.$ The special matrix splitting (2.1) leads us to the subsequent iterative process for solving the three-by-three block SPP given in (1.1): The PESS iterative process. Let $s$ be a positive constant, and let $\Lambda_{1}\in{\mathbb{R}}^{n\times n},\Lambda_{2}\in{\mathbb{R}}^{m\times m}$ $\Lambda_{3}\in{\mathbb{R}}^{p\times p}$ be SPD matrices. For any initial guess ${\bf u}_{0}\in{\mathbb{R}}^{n+m+p}$ and $k=0,1,2\ldots,$ until a specified termination criterion is fulfilled, we compute
@@
where ${\bf u}_{k}=[x_{k}^{T},y_{k}^{T},z_{k}^{T}]^{T},{\bf c}={\mathscr{P}}_{PESS}^{-1}{\bf d},$ and $\mathcal{T}={\mathscr{P}}_{PESS}^{-1}\mathcal{Q}_{PESS}$ is called the iteration matrix for the PESS iterative process. Moreover, the matrix splitting (2.2) introduces a new preconditioner ${\mathscr{P}}_{PESS}$, identified as PESS preconditioner. This preconditioner is a generalization of the previous work in the literature, see for example [9, 27, 30] for specific choices of $\Lambda_{1},\Lambda_{2},\Lambda_{3}$ and $s,$ which are listed in Table 1.
@@
where $r=[r_{1}^{T},r_{2}^{T},r_{3}^{T}]^{T}\in{\mathbb{R}}^{n+m+p}$ and $w=[w_{1}^{T},w_{2}^{T},w_{3}^{T}]^{T}\in{\mathbb{R}}^{n+m+p}.$ Specifying $\hat{X}=\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C$ and $\hat{A}=\Lambda_{1}+sA+s^{2}B^{T}\hat{X}^{-1}B,$ then ${\mathscr{P}}_{PESS}$ can be written in the following way
${\mathscr{P}}_{PESS}=\left[\begin{matrix}I&sB^{T}\hat{X}^{-1}&0\\ 0&I&-sC^{T}\Lambda_{3}^{-1}\\ 0&0&I\end{matrix}\right]\left[\begin{matrix}\hat{A}&0&0\\ 0&\hat{X}&0\\ 0&0&\Lambda_{3}\end{matrix}\right]\left[\begin{matrix}I&0&0\\ -s\hat{X}^{-1}B&I&0\\ 0&s\Lambda_{3}^{-1}C&I\end{matrix}\right].$ (2.13)
@@
1:Solve $(\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C)v_{1}=r_{2}+sC^{T}\Lambda_{3}^{-1}r_{3}$ to find $v_{1};$
@@
3:Solve $(\Lambda_{1}+sA+s^{2}B^{T}(\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C)^{-1}B)w_{1}=v$ to find $w_{1}$
4:Solve $(\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C)v_{2}=sBw_{1}$ to find $v_{2}$;
@@
The implementation of Algorithm 1 requires to solve two linear subsystems with coefficient matrices $\hat{X}=\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C$ and $\hat{A}=\Lambda_{1}+sA+s^{2}B^{T}(\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C)^{-1}B$. Since $\Lambda_{1},\Lambda_{2},\Lambda_{3},\hat{X},$ and $\hat{A}$ are SPD and $s>0$, we can apply the Cholesky factorization to solve these linear subsystems exactly or inexactly by the preconditioned conjugate gradient method. Similarly, we can solve step $6$ of the Algorithm 1.

### 17 May 2024 v2

@@ Abstract
This paper proposes a new parameterized enhanced shift-splitting (PESS) preconditioner to solve the three-by-three block saddle point problem (SPP). Additionally, we introduce a local PESS (LPESS) preconditioner by relaxing the PESS preconditioner. Necessary and sufficient criteria are established for the convergence of the proposed PESS iterative process for any random initial guess. Furthermore, we meticulously investigate the spectral bounds of the PESS and LPESS preconditioned matrices. Moreover, empirical investigations have been performed for the sensitivity analysis of the proposed PESS preconditioner, which unveils its robustness. Numerical experiments are carried out to demonstrate the enhanced efficiency and robustness of the proposed PESS and LPESS preconditioners compared to the existing block diagonal and shift-splitting preconditioners.
Keywords. Krylov subspace iterative process $\cdot$ Preconditioner $\cdot$ Shift-splitting $\cdot$ Saddle point problem $\cdot$ Robust $\cdot$ GMRES
AMS subject classification. 65F08 $\cdot$ 65F10 $\cdot$ 65F50
@@ 1 Introduction
where $\mathcal{A}=\left[\begin{matrix}A&B^{T}&0\\ -B&0&-C^{T}\\ 0&C&0\end{matrix}\right],\,{\bf u}=\left[\begin{matrix}x\\ y\\ z\end{matrix}\right],\,{\bf d}=\left[\begin{matrix}f\\ g\\ h\end{matrix}\right],$ and $A\in{\mathbb{R}}^{n\times n}$ is a symmetric positive definite (SPD) matrix, $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ are the full row rank matrices. Here, $B^{T}$ stands for the transpose of $B,$ and $f\in{\mathbb{R}}^{n},\,g\in{\mathbb{R}}^{m}$ and $h\in{\mathbb{R}}^{p}$ are known vectors. Under the stated assumptions on block matrices A, B, and C, $\mathcal{A}$ is nonsingular, which ensures the existence of a unique solution for the system (1.1).
@@
Due to the growing popularity of three-by-three block SPPs in a variety of applications, it is essential to determine the efficient and robust numerical solution of three-by-three block SPPs (1.1). However, the indefiniteness and poor spectral properties of $\mathcal{A}$ makes challenging to determine the solution of the system (1.1). Furthermore, $\mathcal{A}$ is generally sparse and large; therefore, iterative processes become superior to direct methods. The Krylov subspace iterative process is preferable among other iterative processes investigated widely in the literature due to their minimal storage requirement and feasible implementation [30]. Slow convergence of the Krylov subspace iterative process is a major drawback for system of linear equations of large dimensions. Moreover, the saddle point matrix $\mathcal{A}$ can be very sensitive, which leads to a significant slowdown in the solution algorithm. Therefore, it is important to develop novel, efficient, and robust preconditioners for the fast convergence of the Krylov subspace iterative process, which can handle the sensitivity of the problem (1.1) when the matrix $\mathcal{A}$ is ill-conditioned.
Consider the following partitioning of $\mathcal{A}:$
[…]
@@
then the linear system (1.1) can be identified as the standard two-by-two block SPP [5]. In the last decades, colossal attention has been given to the development of numerical solution methods for standard two-by-two block SPP, such as Uzawa methods [7, 13, 2], Krylov subspace iterative process [17, 26], null space methods [15, 25]. However, the properties of the submatrices in (1.8) are different from the standard two-by-two block SPP. Notice that the leading blocks are skew-symmetric and positive semidefinite in the first partitioning. The second partitioning highlights that the $(1,2)$ block is rank deficient, and the $(2,2)$ block is skew-symmetric. In contrast, in the standard SPP, $(1,2)$ block is full row rank, and the $(2,2)$ block is generally symmetric positive semidefinite. Therefore, the existing literature for solving the two-by-two SPPs (see [5]) may not be applied to solve (1.1) directly, and it is essential to develop new preconditioners for three-by-three block SPP that exploit the specific structure of the saddle point matrix $\mathcal{A}$, which is sensitive in nature.
@@
${\mathscr{P}}_{BD}=\left[\begin{matrix}A&0&0\\ 0&S&0\\ 0&0&CS^{-1}C^{T}\end{matrix}\right]\quad\mbox{and}\quad{\mathscr{P}}_{IBD}=\left[\begin{matrix}\widehat{A}&0&0\\ 0&\widehat{S}&0\\ 0&0&C\widehat{S}^{-1}C^{T}\end{matrix}\right],$ (1.9)
where $S=BA^{-1}B^{T},$ $\widehat{A}$ and $\widehat{S}$ are SPD approximations of $A$ and $S,$ respectively. Although the spectrum of the preconditioned matrices ${\mathscr{P}}^{-1}_{BD}\mathcal{A}$ and ${\mathscr{P}}^{-1}_{IBD}\mathcal{A}$ have good clustering properties, these preconditioners have certain shortcomings, such as they are time-consuming, expensive, require excessive number of iterations, and condition numbers are very large. For more details, see the paper [18]. To overcome these drawbacks of BD preconditioners, researchers have started exploring shift-splitting (SS) preconditioners, which was initially developed for a non-Hermitian system of linear equations by Bai et al. [3] and later for two-by-two block SPPs [10, 23, 11]. Recently, Cao [9] enhanced this idea and introduced the following SS preconditioner ${\mathscr{P}}_{SS}$ and relaxed SS (RSS) preconditioner ${\mathscr{P}}_{RSS}$ for the three-by-three block SPP (1.1)
${\mathscr{P}}_{SS}=\frac{1}{2}\left[\begin{matrix}\alpha I+A&B^{T}&0\\ -B&\alpha I&-C^{T}\\ 0&C&\alpha I\end{matrix}\right]\,\,\mbox{and}\,\,{\mathscr{P}}_{RSS}=\frac{1}{2}\left[\begin{matrix}A&B^{T}&0\\ -B&\alpha I&-C^{T}\\ 0&C&\alpha I\end{matrix}\right],$ (1.10)
where $\alpha>0$ and verified unconditionally convergence of the associated SS iterative process. Here, $I$ stands for the identity matrix of the appropriate dimension. Wang and Zhang [27] generalized the SS preconditioner by introducing a new parameter $\beta>0$ in the $(3,3)$ block. Further, Yin et al. [31] developed an extensive generalized SS (EGSS) preconditioner
@@
for the three-by-three block SPP (1.1), where $\alpha,\beta,\gamma>0$ and $P,Q,W$ are SPD matrices and investigated its convergent properties. There are a number of other preconditioners and iterative processes to solve the three-by-three block SPP (1.1) has been studied recently in the literature [19, 1, 24, 29, 4].
Despite exhibiting favorable performance of SS, RSS and EGSS preconditioners, they do not outperform BD and IBD preconditioners. Therefore, to improve the convergence speed and efficiency of the preconditioners ${\mathscr{P}}_{SS},{\mathscr{P}}_{RSS}$ and ${\mathscr{P}}_{EGSS},$ this paper presents a parameterized enhanced shift-splitting (PESS) preconditioner by introducing a parameter $s>0$ for there-by-three block SPPs. As per the direct correlation between the rate of convergence in Krylov subspace iterative processes and the spectral properties of the preconditioned matrix, we perform in-depth spectral distribution of PESS preconditioned matrix. We summarize the following main contributions of this paper:
  • We propose the PESS iterative process along with its associated PESS preconditioner and its relaxed version known as the local PESS (LPESS) preconditioner to solve the three-by-three block SPP (1.1). The proposed preconditioners generalize the existing SS preconditioners, for example, [9, 27, 31].
@@
  • We empirically show that our proposed preconditioner significantly reduces the condition number of the ill-conditioned saddle point matrix $\mathcal{A}$. Thereby establishing an efficiently solvable, well-conditioned system.
  • Numerical experiments show that the proposed PESS and LPESS preconditioners outperform the existing BD and SS preconditioners. Moreover, we perform sensitivity analysis by perturbing the ill-conditioned matrix $\mathcal{A}$ in (1.1), demonstrating the robustness of the proposed preconditioner.
The structure of this article is as follows. In Section 2, we propose the PESS iterative process and the associated PESS preconditioner. Section 3 is devoted to investigating the convergence of the PESS iterative process. In Section 4, we investigate the spectral distribution of the PESS preconditioned matrix. In Section 5, we present the LPESS preconditioner. In Section 6, we conduct numerical experiments to illustrate the computational efficiency and robustness of the proposed preconditioner. In the end, a brief concluding remark is provided in Section 7.
Notation. The following notations are adopted throughout the paper. For any matrix $A\in{\mathbb{R}}^{n\times n},$ $A^{T},\sigma(A)$ and $\vartheta(A)$ denotes transpose, spectrum and spectral radius of $A,$ respectively. For any $x\in\mathbb{C}^{n},$ $x^{*}$ denotes its conjugate transpose. The symbol $\mathrm{diag}(A,B)$ represents a block diagonal matrix with diagonal blocks $A$ and $B.$ The real and imaginary parts of any complex number $z$ are represented by $\Re(z)$ and $\Im(z)$, respectively. For any symmetric matrix $A\in{\mathbb{R}}^{n\times n},$ $\lambda_{\max}(A)$ and $\lambda_{\min}(A)$ denotes the maximum and minimum eigenvalue of $A,$ respectively. The notation $A>0$ signifies that $A$ is SPD, and $A>B$ identifies $A-B>0.$ Abbreviations used in the paper are as follows:
SPP: saddle point problem, SPD: symmetric positive definite, BD: block diagonal, IBD: inexact block diagonal, SS: shift-splitting, RSS: relaxed shift-splitting, EGSS: extensive generalized shift-spliting, PESS: parameterized enhanced shift-splitting, LPESS: local parameterized enhanced shift-splitting GMRES: generalized minimum residual process.
@@ 2 The parameterized enhanced shift-splitting (PESS) iterative process and the preconditioner
In this section, we proposed a parameterized enhanced shift-splitting (PESS) iterative process and the corresponding PESS preconditioner. Let $s>0$ be a real number. Then, we split the matrix $\mathcal{A}$ in the form
$\mathcal{A}=(\Sigma+s\mathcal{A})-(\Sigma-(1-s)\mathcal{A})=:{\mathscr{P}}_{PESS}-\mathcal{Q}_{PESS},\,\mbox{where}$ (2.1)
${\mathscr{P}}_{PESS}=\left[\begin{matrix}\Lambda_{1}+sA&sB^{T}&0\\ -sB&\Lambda_{2}&-sC^{T}\\ 0&sC&\Lambda_{3}\end{matrix}\right],\mathcal{Q}_{PESS}=\left[\begin{matrix}\Lambda_{1}-(1-s)A&-(1-s)B^{T}&0\\ (1-s)B&\Lambda_{2}&(1-s)C^{T}\\ 0&-(1-s)C&\Lambda_{3}\end{matrix}\right],$
$\Sigma=\mathrm{diag}(\Lambda_{1},\Lambda_{2},\Lambda_{3}),$ and $\Lambda_{1},\Lambda_{2},\Lambda_{3}$ are SPD matrices. The matrix ${\mathscr{P}}_{PESS}$ is nonsingular for $s>0.$ The special matrix splitting (2.1) leads us to the subsequent iterative process for solving the three-by-three block SPP given in (1.1):
The PESS iterative process. Let $s$ be a positive constant, and let $\Lambda_{1}\in{\mathbb{R}}^{n\times n},\Lambda_{2}\in{\mathbb{R}}^{m\times m}$ $\Lambda_{3}\in{\mathbb{R}}^{p\times p}$ be SPD matrices. For any initial guess ${\bf u}_{0}\in{\mathbb{R}}^{n+m+p}$ and $k=0,1,2\ldots,$ until a specified termination criterion is fulfilled, we compute
@@
where ${\bf u}_{k}=[x_{k}^{T},y_{k}^{T},z_{k}^{T}]^{T},{\bf c}={\mathscr{P}}_{PESS}^{-1}{\bf d},$ and $\mathcal{T}={\mathscr{P}}_{PESS}^{-1}\mathcal{Q}_{PESS}$ is called the iteration matrix for the PESS iterative process.
Moreover, the matrix splitting (2.2) introduces a new preconditioner ${\mathscr{P}}_{PESS}$, identified as PESS preconditioner. This preconditioner is a generalization of the previous work in the literature, see for example [9, 27, 31] for specific choices of $\Lambda_{1},\Lambda_{2},\Lambda_{3}$ and $s,$ which are listed in Table 1.
@@
where $r=[r_{1}^{T},r_{2}^{T},r_{3}^{T}]^{T}\in{\mathbb{R}}^{n+m+p}$ and $w=[w_{1}^{T},w_{2}^{T},w_{3}^{T}]^{T}\in{\mathbb{R}}^{n+m+p}.$ Specifying $\widehat{X}=\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C$ and $\widehat{A}=\Lambda_{1}+sA+s^{2}B^{T}\widehat{X}^{-1}B,$ then ${\mathscr{P}}_{PESS}$ can be written in the following way
${\mathscr{P}}_{PESS}=\left[\begin{matrix}I&sB^{T}\widehat{X}^{-1}&0\\ 0&I&-sC^{T}\Lambda_{3}^{-1}\\ 0&0&I\end{matrix}\right]\left[\begin{matrix}\widehat{A}&0&0\\ 0&\widehat{X}&0\\ 0&0&\Lambda_{3}\end{matrix}\right]\left[\begin{matrix}I&0&0\\ -s\widehat{X}^{-1}B&I&0\\ 0&s\Lambda_{3}^{-1}C&I\end{matrix}\right].$ (2.4)
@@
1:Solve $\widehat{X}v_{1}=r_{2}+sC^{T}\Lambda_{3}^{-1}r_{3}$ to find $v_{1};$
@@
3:Solve $\widehat{A}w_{1}=v$ to find $w_{1}$
4:Solve $\widehat{X}v_{2}=sBw_{1}$ to find $v_{2}$;
@@
The implementation of Algorithm 1 requires to solve two linear subsystems with coefficient matrices $\widehat{X}$ and $\widehat{A}$. Since $\Lambda_{1},\Lambda_{2},\Lambda_{3},\widehat{X},$ and $\widehat{A}$ are SPD and $s>0$, we can apply the Cholesky factorization to solve these linear subsystems exactly or inexactly by the preconditioned conjugate gradient method.